OpenStudy (anonymous):

Find an equation of the tangent to the curve y = e^x that passes thru the origin.

OpenStudy (anonymous):

So first we use the derivative to find the slope. Can you do that?

OpenStudy (anonymous):

Now we know \[\Large m=\frac{y_2-y_1}{x_2-x_1}\] So \[\Large \frac{dy}{dx}e^x=\frac{e^x-0}{x- 0}\]

OpenStudy (anonymous):

is that the answer?

OpenStudy (anonymous):

No, but using this equation we can solve for \(x\), to find which point to use to get the slope.

OpenStudy (anonymous):

Then we can put the line in the form \( y = mx +b \) and we know \(b=0\) since the origin is the intercept.

OpenStudy (anonymous):

so the answer is e^x

OpenStudy (anonymous):

No, \(\large e^x\) isn't even a line! You need to solve for \(x\) in \[\Large \frac{dy}{dx}e^x=\frac{e^x-0}{x- 0}\]

OpenStudy (anonymous):

First start of by finding the derivative.

OpenStudy (anonymous):

the derivative is e^x

OpenStudy (anonymous):

and then e^0 is 1

OpenStudy (anonymous):

Okay so when does \[\large e^x = \frac{e^x}{x}\] What must x be for both sides to be equal?

OpenStudy (anonymous):

1

OpenStudy (anonymous):

Good, so now we can use it to find \(m\) by plugging \(1\) into \(\large e^x\).

OpenStudy (anonymous):

And our answer will be \(y = mx+b\). We know \(b = 0\) and \(m = e^1\).

OpenStudy (anonymous):

So the line is \(y=e \cdot x\). Make sense?

OpenStudy (anonymous):

yes...Thank You

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=e%5Ex%2Cex And this confirms it!

OpenStudy (anonymous):

Thank you ....Have a nice day :)